OpenStudy (anonymous):

how do i find the sixth root of 729w^729?

OpenStudy (anonymous):

3^6=729...it's a start

OpenStudy (anonymous):

thats the answer?

OpenStudy (zarkon):

no...that is a hint

OpenStudy (anonymous):

would you do 729/ 1/6?

OpenStudy (zarkon):

\[\sqrt[6]{729w^{729}}\] \[=\sqrt[6]{729}\sqrt[6]{w^{729}}\]=...

OpenStudy (anonymous):

got it its 3w^121 6radicalw^3